Does the ring Z7[x]/<x^2+3> have nilpotent elements? justify.
1
Expert's answer
2020-10-12T18:23:33-0400

Using the necessary condition for nilpotent elements - they cannot be invertible, we fix the element of the ring ax+b and try to solve this equation:

"(ax+b)(cx+d)=1\\
acx^2+(ad+bc)x+bd=1\\
(ad+bc)x+(bd-3ac)=1\\
\begin{cases}
 ad+bc=0\\
 bd-3ac=1
 \end{cases}\\
Because\ \mathbb{Z}_7 \ is \ a \ field\ from \ first\ equation:\\
c=-\frac{ad}{b}\\
Put\ in\ second:\\
bd+\frac{3a^2d}{b}=1"

It is easy to see that this equation is solvable for b not equal to 0

If b=0:

"ax(cx+d)=1\\
adx-3ac=1"

And this equation is easily solvable for a not equal to 0, which means that there are no nilpotent elements in the ring